OpenStudy (anonymous):

When solving a radical equation, Beth and Kelly came to two different conclusions. Beth found a solution, while Kelly's solution did not work in the equation. Create and justify two situations: one situation where Beth is correct and a separate situation where Kelly is correct.

OpenStudy (amoodarya):

\[\sqrt{x}-x=0\] \[\sqrt{x}=x\] \[(\sqrt{x}=x)^2\] \[x=x^2 \rightarrow x-x^2=0 \rightarrow x(1-x)=0 \rightarrow x=0 ,1\] both of roots are correct

OpenStudy (amoodarya):

another situation \[\sqrt{x}+x=0\] \[\sqrt{x}=-x\] \[(\sqrt{x}=-x)^2\] \[x=x^2 \rightarrow x(1-x)=0\] the same equation so x=0,1 x=0 is true (check) but x=1 is not true (check) \[\sqrt{1}+1\neq0\]